OpenStudy (anonymous):

If C is 24 and BAG is 71 then DEAF is: a. 90 b. 92 c. 93 d. 94

OpenStudy (anonymous):

Is this one of the weird question from sideways arithmetic from waysides shcool

OpenStudy (anonymous):

i dont know

OpenStudy (shubhamsrg):

considering all this an AP we see a+2d =24 and a+ a+d +a +6d=71 =>3a + 7d =71 =>a=26 d=-1 soDEAF = a+3d + a+4d + a + a+5d = 4a + 12d =92 ! voila ! :D

OpenStudy (shubhamsrg):

i was just writing it,,i never expected my ans to actually match! B|

OpenStudy (anonymous):

It seems correct by the way nice logic applied

OpenStudy (shubhamsrg):

hmmm.. :)